What pressure is exerted by a gas at 25oC at a volume of 0.044 L and 0.00685 moles of gas are present? Volume = _____________
1
Expert's answer
2020-05-20T13:44:39-0400

PV=nRT

P=nRT/V

P=0.00685*8.314*298/0.044= 385.7 kPa